# ESDD2: Environment-Aware Speech and Sound Deepfake Detection Challenge Evaluation Plan

### v1

@@ III Baseline
TABLE IV: Performance comparison on the validation, evaluation, and test sets.
Original EER Speech EER Env EER F1-score
Val set 0.0031 0.0172 0.3766 0.9462
@@ IV Evaluation Criteria
In addition, we introduce three auxiliary metrics based on equal error rate (EER): $\mathrm{EER}_{\text{original}}$, $\mathrm{EER}_{\text{speech}}$ and $\mathrm{EER}_{\text{env}}$. $\mathrm{EER}_{\text{original}}$ is the EER for distinguishes the Original class from all other classes. Similarly, $\mathrm{EER}_{\text{speech}}$ is the EER for detecting spoofed speech component, while $\mathrm{EER}_{\text{env}}$ is the EER for detecting spoofed environmental component. The three EER metrics are included for diagnostic and analysis purposes only; they will not be considered for leaderboard ranking.
@@ V Submission Guideline
The challenge has two phases: the preparation phase, where participants submit evaluation set predictions and the leaderboard reflects evaluation set rankings; and the final evaluation phase, where participants submit test set predictions and the leaderboard determines the final ranking. Each team may submit up to ten results in the final evaluation phase, with the best three used for final ranking to ensure fairness. After the competition, we will release the metadata of the evaluation and test sets to support reproducibility and future research.

### v2

@@ III Baseline
TABLE IV: Performance (EER % and F1-score) of baseline model on the validation, evaluation, and test sets of CompSpoofV2.
Original EER Speech EER Env. EER F1-score
Val. set 0.0031 0.0172 0.3766 0.9462
@@ IV Evaluation Criteria
In addition, we introduce three auxiliary metrics based on equal error rate (EER): $\mathrm{EER}_{\text{original}}$, $\mathrm{EER}_{\text{speech}}$ and $\mathrm{EER}_{\text{env}}$. $\mathrm{EER}_{\text{original}}$ is the EER for distinguishes the Original ($i=0$) class from all other four classes ($i=\{1,2,3,4\}$. Similarly, $\mathrm{EER}_{\text{speech}}$ is the EER for detecting spoofed speech component, while $\mathrm{EER}_{\text{env}}$ is the EER for detecting spoofed environmental component. The three EER metrics are included for diagnostic and analysis purposes only; they will not be considered for leaderboard ranking.
@@ V Submission Guideline
The challenge has two phases: the preparation phase and the final phase. During the preparation phase, participants submit evaluation set predictions and the leaderboard reflects evaluation set rankings; and the final evaluation phase, where participants submit test set predictions and the leaderboard determines the final ranking. Each team may submit up to ten results in the final evaluation phase, with the best three used for final ranking to ensure fairness. After the competition, we will release the metadata of the evaluation and test sets to support reproducibility and future research.